OpenStudy (anonymous):

Find the limit, if it exists. If it does not exist, enter DNE. lim x^3 − 2x + 3/ x-> ∞ 4 − 5x^2 I answered ∞, but it was wrong. Please explain how to get to the answer, thank you.

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ \lim_{x\to\infty} \dfrac{x^3-2x+3}{4-5x^2}\hspace{.33em}\\~\\ \normalsize \text{applying L'hopital's rule}\hspace{.33em}\\~\\ =\lim_{x\to\infty} \dfrac{\dfrac{d}{dx}\left(x^3-2x+3\right)}{\dfrac{d}{dx}\left(4-5x^2\right)}\hspace{.33em}\\~\\ =\lim_{x\to\infty} \dfrac{\left(3x^2-2\right)}{\left(-10x\right)}\hspace{.33em}\\~\\ \normalsize \text{applying L'hopital's rule again} \hspace{.33em}\\~\\ =\lim_{x\to\infty} \dfrac{\dfrac{d}{dx}\left(3x^2-2\right)}{\dfrac{d}{dx}\left(-10x\right)}\hspace{.33em}\\~\\ =\lim_{x\to\infty} \dfrac{6x}{-10}\hspace{.33em}\\~\\ =-\dfrac{3\times \infty }{5}\hspace{.33em}\\~\\ =-\infty }\end{align}\)

OpenStudy (mathmath333):

but m not sure if this solution is totally correct